March 8, 2017, Production & Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Wednesday March 8, 2017 at 2:15 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326-327.

Meeting Connections

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

Reconstruction & Analysis Studies Status


  1. Announcements
  2. Data Production
  3. Analysis Focus
  4. This Week's Studies
  5. Any other studies
  6. Upcoming Study Schedule

This Week's Topics

  • Talks: Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • Talks should be limited to 10 + 5 minutes.

Data Production

  1. Monitoring Update --- Thomas Britton
  2. Calibration Update --- Sean Dobbs
  3. Processing Update --- Alex Austregesilo

Analysis Focus: Total Cross Sections

  1. J/Ψ, φ --- Lubomir
  2. η, ρ, ω, φ --- Simon
  3. φ --- Thomas
  4. Any others?

Studies: Hit Efficiencies

(i.e. Is the detector working?) Updates on experiment, simulation, and the comparison between the two.

  1. CDC Hit Efficiencies --- Naomi Jarvis
  2. FDC Hit Efficiencies --- Alex Austregesilo
  3. BCAL Hit Efficiencies --- Elton Smith
  4. FCAL Hit Efficiencies --- Jon Zarling

Upcoming Study Schedule

  • Updates on experiment, simulation, and the comparison between the two.
  • Summary/overviews only. Detailed discussions should be held in the appropriate working groups.
  • All talks should be limited to 10 + 5 minutes.
  • Next Week: Alignment & Track/Shower Efficiencies (i.e. Is the reconstruction working?)
  1. Drift chamber alignment --- Mike Staib
  2. Tracking Efficiencies --- Paul Mattione
  3. BCAL Shower Efficiencies --- BCAL Group
  4. FCAL Shower Efficiencies --- FCAL Group
  • Week after next: Other (Any of below if update ready)
  1. Track / Hit Matching: BCAL, FCAL, TOF, SC
  2. Efficiencies: TOF, SC
  3. Means & Resolutions (time, energy, dE/dx): Tracking, BCAL, FCAL, SC, TOF
  4. Uncertainties: PID (BCAL, FCAL, TOF), Kinfit (BCAL, FCAL, tracking)
  5. Channel/Analysis Studies: Branching ratios, cross sections, SDMEs, beam asymmetries
  6. Other reconstruction/analysis issues
  • The following week: Beamline & Triggering
  1. Flux --- Justin
  2. Beam energy --- Beamline Group
  3. Polarization (TPOL & lineshape) --- Beamline Group
  4. Trigger emulation --- Alex Somov
  5. Triggering efficiency --- Alex Somov


Data Production

  • Monitoring Update --- The high-level monitoring plots are generally looking good, except for the tagger calibrations.
  • Calibration Update --- Made improvement on channel-to-channel offsets, moved calibrations into default variation over the weekend. Still refining the tagger and calibrations, as well as the calibrations for high rate running.
  • Processing Update --- The latest monitoring launch was started this morning. The jobs usually run overnight, and the results are feeding into the calibration work.

Analysis Focus

  • Dilepton Working Group had an organizing meeting, next meeting scheduled for next Thursday. Currently looking at E/p cuts and reconstruction performance. Also looking into phi cross sections in various decay modes. Lubomir: background in phi->e+e- is still too high without the kinematic fit CL cut.
  • Simon started looking at the Spring 2017 data, still using the PS acceptance from Spring 2016. Limited by tagger calibrations
  • Thomas is looking at phi -> K+K-

This Week's Studies

  • CDC Hit Efficiencies
    • Study of CDC efficiency vs. distance from wire, efficiency is flat up to 3mm away, ~97%. Looking at HV scans, the efficiency for all tracks varies by just ~2%, while for tracks at theta ~ 90%, i.e. those completely contained in the CDC, the efficiency varies by ~10%.
    • Suggestions for improvement include explicitly taking into account detector and looking at tracks from exclusive reactions.
    • Dead straws marked in CCDB for data up to run 30421, need to make sure they get into MC
  • FDC Hit Efficiencies
    • Pseudoefficiency ~90% in recent runs (see Run meeting)
  • BCAL Hit Efficiencies
    • Nothing New
  • FCAL Hit Efficiencies
    • Single channels are missing, these vary over the run. Working on extracting the information using LED pulsers for the current run.
    • Mismatch on inner rings information due to missing trigger simulation, and perhaps other efficiency issues